### v3

# Time Evidence Fusion Network: Multi-source View in Long-Term Time Series Forecasting

Abstract
In practical scenarios, time series forecasting necessitates not only accuracy but also efficiency. Consequently, the exploration of model architectures remains a perennially trending topic in research. To address these challenges, we propose a novel backbone architecture named Time Evidence Fusion Network (TEFN) from the perspective of information fusion. Specifically, we introduce the Basic Probability Assignment (BPA) Module based on evidence theory to capture the uncertainty of multivariate time series data from both channel and time dimensions. Additionally, we develop a novel multi-source information fusion method to effectively integrate the two distinct dimensions from BPA output, leading to improved forecasting accuracy. Lastly, we conduct extensive experiments to demonstrate that TEFN achieves performance comparable to state-of-the-art methods while maintaining significantly lower complexity and reduced training time. Also, our experiments show that TEFN exhibits high robustness, with minimal error fluctuations during hyperparameter selection. Furthermore, due to the fact that BPA is derived from fuzzy theory, TEFN offers a high degree of interpretability. Therefore, the proposed TEFN balances accuracy, efficiency, stability, and interpretability, making it a desirable solution for time series forecasting.
Index Terms:
Time Series, Forecasting, Evidence Theory, Information Fusion, Neural Network
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
The evolution of various phenomena is inherently tied to the progression of time, leading to an increasing prevalence of time series data in a multitude of application domains [1], such as medical diagnosis [2, 3, 4], education [5], weather [6, 7], power system [8], and so on [9, 10, 11]. The importance of time series forecasting cannot be overstated. There are also many methods for time series forecasting. The initial methods came from statistics, such as Autoregressive (AR) model, Moving Average (MA) model, Autoregressive Moving Average (ARMA) model, and Autoregressive integrated moving average (ARIMA) model [12]. Deep learning is a powerful method with wide applications in various disciplines [13, 14]. In recent years, time series prediction models based on neural networks have become a research hotspot. These neural networks have evolved from the simplest deep neural networks (DNN) [15]. Until now, there are already many backbones that can be used for time series, such as Recurrent Neural Network (RNN) [16], Long Short-Term Memory (LSTM) [17], Transformer [18, 19, 20, 21], and so on. Large Language Models (LLM) for time series also perform well in cross dataset forecasting tasks [22, 23, 24].
There is uncertainty in time series data, and modeling uncertainty on time series is effective. Evidence theory is an uncertain reasoning framework based on mass function [25, 26, 27, 28]. Previous studies have shown that evidence theory can effectively utilize the uncertainty of time-series data and achieve positive application results [29, 30]. On the other hand, evidence theory is also a method of information fusion that can effectively aggregate different information sources and make effective decisions [31, 32]. Meanwhile, evidence theory can also be applied in neural networks to accomplish tasks such as classification [33]. Therefore, in order to reduce the error of time series forecasting, we propose the Time Evidence Fusion Network (TEFN) from the perspectives of evidence theory and information fusion. TEFN is a lightweight neural network that considers the time and channel dimensions of multivariate time series as different sources of information in Figure 1.
Figure 1: The channel dimension and time dimension of multivariate time series can be considered as different sources of information. Firstly, merge information sources of the same type, and finally integrate information sources from both channel and time dimensions to complete the final prediction.
In TEFN, we propose a novel Basic Probability Assignment (BPA) Module. Based on the evidence theory of decision-making process, BPA maps different sources of information to distributions related to the target outcome. The target distribution is achieved by the fuzzy membership of different fuzzy sets [34]. In image processing, convolution condenses the required features of the target [35]. Unlike convolution, BPA is an expansion process in Figure 2. On the contrary, BPA considers all possibilities in the event space composed of subsets of the sample space, expanding on different events, which is more effective for simple structures like time series and can fully explore the different hidden information inside. For different distributions generated by multiple information sources, we consider sources as samples and use the expected forecasting values as the final result.
Figure 2: Comparison between convolution and basic probability assignment: convolution condenses information while basic probability assignment diverges to consider different possibilities.
In this article, we focus on the practical application details of TEFN, including its accuracy, memory usage, training speed, parameter sensitivity, and interpretability. At the same time, in order to ensure fairness, the experimental data and configuration files for comparison are all from open-source projects. TEFN achieves state-of-the-art (SOTA) on multiple real datasets, while training time and model parameters are much smaller than models based on Transformer [20, 36]. For random hyperparameter selection, TEFN exhibits weak fluctuations and is a stable model. The contributions of this article are as follows:
  • A novel evidential backbone BPA designed for extracting potential distributions from simple data structures is introduced.
  • A BPA-based neural network, TEFN, is introduced for time series forecasting, outpacing prior research in terms of accuracy, efficiency, stability, and interpretability.
  • The manuscript elucidates the incorporation of evidence theory concepts, such as the mass function, into the architecture of neural networks, offering a novel approach to uncertainty reasoning in machine learning.
This paper is organized as follows. We begin by reviewing relevant literature in the Section 2. Next, in the Section 3, we delve into the architecture and underlying mathematical principles of TEFN. In the Section 4, we present a series of experiments that demonstrate TEFN’s superior performance. Finally, we conclude the paper by summarizing our key findings in the Section 5.
2 Related Work
2.1 Dempster-Shafer Theory
Dempster-Shafer theory, also known as Evidence Theory, operates on a different framework compared to traditional probability theory [25, 26]. It relies on the concept of mass functions, which provide a more flexible representation of uncertainty, accommodating weaker constraints than traditional probabilities. All independent and mutually exclusive sample sets that exist in the system are called frames of discrete (FOD), represented as $\Omega$.
When there are multiple sources of information, information fusion is usually required to make decisions. Fusing two mass functions, denoted as $m_{1}$ and $m_{2}$, the fused mass function $m$ is determined using the Dempster-Shafer Rule (DSR), as described in Equation (1). The symbol variables $A$, $B$, and $C$ are focal elements, which are typically combinations of all independent samples in the FOD, which are often composed of target labels, such as the classification targets.
$m(A)=\sum_{B\cap C=A}m_{1}(B)*m_{2}(C)$ (1)
2.2 Basic Probability Assignment
BPA is a core concept in evidence theory, used to map real data $x$ into mass distributions. Mass distributions describe the degree of support for various hypotheses in the event space, providing a basis for reasoning and decision-making. The generation of BPA $m$ is often based on fuzzy logic, where fuzzy membership functions $\mu$ are used to map data points to different fuzzy sets, thereby constructing the target distribution in Equation (2) where $w$ and $b$ is weight and bias of the membership function. BPA can also be generated by assigning support degrees through Gaussian distributions $N$ (probability density function $f$) where $\mu$ and $\sigma$ are expectation and variance in Equation (3).
$\begin{split}m=\mu(x)=wx+b\end{split}$ (2)
$\begin{split}&x\sim N(\mu,\sigma)\\ &m=f(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}\\ \end{split}$ (3)
The advantage of BPA lies in its ability to effectively handle uncertainty and improve the accuracy and reliability of predictions by fusing multi-source information.
2.3 Evidence Decision Making
The process of Evidence Decision Making can be delineated into several sequential steps: basic probability assignment (BPA), evidence fusion, and decision making (DM) [34, 37]. BPA serves as a pivotal initial phase, constituting a mapping from raw data to mass distribution, akin to the path from raw data to probability distribution. Subsequently, evidence fusion integrates distinct mass distributions from various data channels, culminating in a rationalized mass distribution. The most classic fusion method is the DSR [25, 26] mentioned in Section 2.1. Finally, decision making transpires as the progression from the attained distribution towards the designated target. The most classic decision-making method is to transform the mass distribution into the probability distribution of the target, called pignistic probability transformation (PPT) [38] in Equation (4), where $x$ is a target label and $e$ are combinations of all possible target labels.
$p(x)=\sum_{x\in e}\frac{m(e)}{(1-m(\emptyset)|e|)}$ (4)
2.4 Linear Time Series Model
There are many linear time series forecasting models (LTSFM) [12, 39, 40, 41]. LTSFM offer several advantages for modeling and forecasting time series data:
  • The simplicity and transparency make them easy to implement, interpret, and analyze.
  • This leads to faster training and inference times, making them suitable for real-time applications and large datasets.
  • No additional encoding layers are required.
A classic LTSFM model is Dlinear [39]. DLinear serves as a simple yet effective baseline for long-term time series forecasting. It is a model that employs a seasonal-trend decomposition technique to separate the trend and seasonal components of the time series data. Each component is then independently modeled using a single-layer linear regression. The final prediction is obtained by summing the outputs of the two linear regressions.
TEFN shares similarities with Dlinear in that both models utilize linear regression for prediction. However, there are key differences:
  • DLinear employs a simpler decomposition method compared to TEFN, which may limit its ability to capture intricate temporal dynamics.
  • DLinear focuses on capturing linear relationships and trends in the data, while TEFN focuses on the mining of internal hidden relationships, so it strengthens predictions through dimensionality increase and uncertain reasoning.
  • TEFN is a general framework that does not require identification and prediction of trends and periodicity, while Dlinear is more like a feature modeling prediction of time series.
3 Time Evidence Fusion Network
The structure of TEFN is inherently straightforward, achieving remarkable performance with a design featuring only two BPA modules, akin to employing just two convolutional kernels. We posit that the novel architecture of TEFN embodies the essence of a pure backbone model, stripped of domain-specific intricacies. This minimalist approach allows TEFN to focus solely on capturing the fundamental temporal patterns within the data, making it a versatile and adaptable foundation for various time series analysis tasks. The overall structure of TEFN is depicted in Figure 3. Next, we will elaborate on each module one by one.
Figure 3: The overall structure of TEFN: Establish corresponding mass functions for the time dimension and channel dimension as the target dimensions of BPA.
Figure 4: The operation of a BPA module: weight matrix and bias matrix are both learnable parameters. BPA elevates the dimensionality of data by adding a dimension to distinguish different events.
3.1 Time Normalization and De-normalization
To expedite convergence towards local optima and enhance training efficiency, TEFN incorporates a normalization technique inspired by the Stationary model [42]. This process computes the mean $\mu$ as shown in Equation (5) and variance $\sigma^{2}$ as shown in Equation (6) for each segment of the time series $x$. The subsequent normalization and de-normalization operations, presented in Equation (7) and Equation (8), where $\hat{y}$ is forecasting result, respectively, act as symmetrical transformations that standardize the time series data. The neural network operation $Net(\cdot)$ is then applied to the normalized time series $x_{Norm}$, ensuring that the model learns from data with consistent scale and distribution. This normalization strategy effectively mitigates the influence of outliers and initial data scale, fostering faster convergence and improved stability during training.
$\mu=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}$ (5)
$\sigma^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}(x_{i}-\mu)^{2}=\frac{1}{|x|}\sum_{x_{i}\in x}x_{i}^{2}-\mu^{2}$ (6)
$x_{Norm}=\frac{x-\mu}{\sigma}$ (7)
$\hat{y}=\sigma*\hat{y}_{Norm}+\mu=\sigma*Net(x_{Norm})+\mu$ (8)
3.2 Time Dimension Projection
In TEFN, we employ a linear layer $project(\cdot)$ to project dimensionality to an expansion dimensionality, transforming the input time series $x$ with a length of $L_{in}$ into a sequence $x^{\prime}$ with a length of $L_{in}+L_{pred}$ in Equation (9). Here, $L_{pred}$ represents the desired length of the prediction horizon. This operation is analogous to the time projection mechanism utilized in other neural network architectures, such as TimesNet [43].
$x^{\prime}=project(x)=x\times W_{p}+b$ (9)
By extending the input vector to match the prediction length, the time projection layer facilitates the direct generation of future values based on the learned temporal patterns. This approach simplifies the model design and enables efficient prediction without the need for iterative or recursive computations, as commonly observed in autoregressive models.
3.3 Basic Probability Assignment Module
In TEFN, we leverage the principles of Evidence Theory to model the uncertainty and ambiguity inherent in time series data. According to Evidence Theory, each dimension of the time series can be represented as a mass function within a common event space. This event space is typically defined as the power set $2^{S}$ of the sample space $S$, where $S$ represents the possible values or labels for the data. Since time series forecasting is a regression problem, the size of the sample space needs to be used as a hyperparameter.
This representation allows us to capture the fuzzy features of the input time series, where each dimension generates a mass distribution within the event space. To achieve this, we utilize parameterized fuzzy membership functions, as defined in Equation (10). Here, $D$ represents the dimensionality of the time series, encompassing both the time dimension $T$ and the channel dimension $C$. The indices $i$ and $j$ correspond to the specific dimensions within $D$, while $k$ represents the index of the event space dimension, denoted as $F$. The fuzzy membership function $\mu_{k}$ is defined in Equation (10) where $w_{D,j,k}$ and $b_{D,j,k}$ are the slope and intercept parameters of the membership function. These parameters are learned during training and allow the model to capture the varying degrees of membership for each data point within the event space. Figure 4 shows the visible process of BPA module.
$m_{D,i,j,k}=\mu(x_{Norm,i,j})=w_{D,j,k}*x_{Norm,i,j}+b_{D,j,k}$ (10)
By employing evidence theory and fuzzy membership functions, TEFN effectively models the uncertainty in time series data and enables the generation of robust and accurate predictions. This approach provides a principled way to handle the inherent fuzziness and ambiguity in real-world time series, leading to improved performance and reliability in various applications.
To introduce non-linear characteristics into TEFN and enhance its modeling capabilities, we can modify the fuzzy membership function used in the BPA module. Instead of the linear form presented in Equation (10), we can adopt non-linear functions that capture more complex relationships within the data. Several options exist for non-linear membership functions, including:
  • Gaussian Function: This function models the data distribution as a bell-shaped curve such as Equation (3), capturing the likelihood of each data point belonging to different membership levels. The parameters of the Gaussian function can be learned to adapt to the specific characteristics of the time series data.
  • Sigmoid Function: This function smoothly transitions between 0 and 1, providing a flexible way to represent the membership of data points to different levels in Equation (11). The sigmoid function can capture both gradual and abrupt changes in membership values.
    $\sigma(x)=\begin{cases}0&\text{if }x<0\\ \frac{1}{1+e^{-x}}&\text{if }x\geq 0\end{cases}$ (11)
  • Piecewise Functions: These functions consist of multiple linear or non-linear segments, allowing for more precise control over the shape of the membership function in Equation (12) where $g(x)$ is a non-linear function. This approach is useful for capturing complex patterns and relationships within the data.
    $f(x)=\begin{cases}a_{1}x+b_{1}&\text{if }c_{1}x+d_{1}\leq 0\\ g_{1}(x)&\text{if }c_{2}x+d_{2}>0\text{ and }c_{3}x+d_{3}\leq 0\\ a_{2}x+b_{2}&\text{if }c_{4}x+d_{4}>0\\ \vdots&\\ a_{n}x+b_{n}&\text{otherwise}\end{cases}$ (12)
By utilizing non-linear membership functions, TEFN can effectively model the non-linear dynamics and interactions within the time series data. This enhancement enables TEFN to capture more subtle patterns and relationships, leading to improved accuracy and generalization in various time series forecasting tasks.
3.4 Expectation Fusion
TEFN addresses the challenge of handling multivariate time series data with two distinct dimensions: the time dimension $T$ and the channel dimension $C$. This results in the paralleled generation of different mass distributions for each dimension, denoted as $m_{C}$ and $m_{T}$. To integrate these mass distributions and produce a unified prediction, TEFN employs a fusion method based on expectations, as defined in Equation (13). Here, $y_{j,k}$ represents the fusion parameters, and $m^{\prime}_{i,j,k}$ corresponds to the transformed mass function for each dimension.
$\begin{split}\hat{y}_{Norm,D,i,j}=E_{D,F}(y_{j,k})&=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}y_{j,k}*m^{\prime}_{i,j,k}\\ &=\sum_{\begin{subarray}{c}j\in[1,|D|]\\ k\in[1,|2^{S}|]\end{subarray}}m_{i,j,k}\end{split}$ (13)
The fusion operation in Equation (13) can be interpreted as a linear transformation of the mass functions, where the summation of $y_{j,k}*m^{\prime}_{i,j,k}$ effectively combines the information from different dimensions. The expectation is then calculated by summing the transformed mass functions, resulting in a single mass function representing the fused prediction.
TEFN deliberately avoids using the DSR for several reasons:
  1. 1.
    The computational complexity of DSR is significantly higher compared to the expectation-based fusion method. Additionally, DSR involves point-to-point multiplication of mass distributions within the same dimension, which can lead to accuracy loss when dealing with high-dimensional data due to finite precision limitations.
  2. 2.
    DSR is sensitive to extreme distributions, particularly those with a single element having a value of 1. This sensitivity can result in the final fused result being dominated by that single distribution, potentially leading to conflicts and loss of information from other distributions.
By employing the expectation-based fusion method, TEFN achieves a balance between computational efficiency and accuracy. This approach effectively integrates information from different dimensions while mitigating the risks of conflicts and accuracy loss associated with DSR. The final normalized prediction results are then aggregated across dimensions using the equation in Equation (14) and restored to their original scale through the de-normalization layer.
$\hat{y}_{norm}=\sum_{D\in\{T,C\}}\hat{y}_{Norm,D}=\hat{y}_{Norm,T}+\hat{y}_{Norm,C}$ (14)
4 Experiments
The experimental section aims to comprehensively evaluate TEFN’s performance in long-term time series forecasting. Utilizing diverse datasets, TEFN will be benchmarked against established models, measuring accuracy, efficiency, and robustness. We will analyze its performance across different prediction horizons, assess its sensitivity to hyperparameters, and conduct ablation studies to understand the impact of individual components. Lastly, we will explore TEFN’s interpretability by analyzing the basic probability assignments generated by its BPA module.
4.1 Dataset
Our research employs the TEFN for the challenging task of long-term time series prediction. To ensure the model’s effectiveness and robustness across diverse scenarios, we meticulously selected five representative and extensively used datasets. These datasets encompass a wide range of real-world applications, providing a comprehensive evaluation of TEFN’s performance. The selected datasets are:
  • Electricity [44]: This dataset contains electricity load diagrams for a power grid from 2011 to 2014, offering insights into energy consumption patterns over time. It presents a challenging forecasting task due to its seasonal variations and potential anomalies.
  • ETT (4 subsets) [19]: This dataset comprises four subsets with varying lengths and complexities, representing diverse time series forecasting challenges. It allows us to assess TEFN’s ability to handle different data characteristics and prediction horizons.
  • Exchange [45]: This dataset captures exchange rate fluctuations between major currencies. It presents a dynamic and volatile environment, testing TEFN’s ability to capture complex market trends and predict future exchange rates accurately.
  • Traffic [46]: This dataset contains traffic flow data collected from sensors in a city. It reflects the temporal and spatial patterns of traffic congestion, providing a challenging task for forecasting traffic flow and congestion levels.
  • Weather [47]: This dataset includes historical weather data, encompassing various meteorological parameters. It allows us to evaluate TEFN’s performance in predicting weather patterns, which are crucial for numerous applications like agriculture, disaster management, and transportation planning.
Each of these datasets is characterized by its large size and diverse features, offering a rich and challenging environment for evaluating TEFN’s capabilities. The extensive nature of the data ensures that the model is tested on a wide range of scenarios, providing valuable insights into its performance and generalizability. More dataset features are in the Table I.
TABLE I: The quantitative features of the dataset
Dataset Dimension Dataset Size
Electricity 321 (18317, 2633, 5261)
ETTm1, ETTm2 7 (34465, 11521, 11521)
ETTh1, ETTh2 7 (8545, 2881, 2881)
Traffic 862 (12185, 1757, 3509)
Weather 21 (36792, 5271, 10540)
Exchange 8 (5120, 665, 1422)
TABLE II: This table presents a comprehensive comparison of prediction errors across various models for long-term time series forecasting tasks. The models are evaluated on four different prediction lengths: $L_{pred}\in\{96,192,336,720\}$, providing a thorough assessment of their performance under varying forecasting horizons. The color coding in the table highlights the ranking of each model’s performance across all datasets and prediction lengths, allowing for easy identification of the top-performing models.
Models
TEFN
iTransformer
Rlinear
PatchTST
Crossformer
TiDE
TimesNet
ETSformer
LightTS
Dlinear
FEDformer
Stationary
Autoformer
[48]
[49]
[36]
[20]
[50]
[43]
[51]
[40]
[39]
[52]
[42]
[45]
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.343
0.367 0.334 0.368
0.355
0.376
0.329 0.367
0.404
0.426
0.364
0.387
0.338
0.375
0.375
0.398
0.374
0.400
0.345
0.372
0.379
0.419
0.386
0.398
0.505
0.475
192
0.381
0.383
0.377
0.391
0.391
0.392
0.367 0.385
0.450
0.451
0.398
0.404
0.374
0.387
0.408
0.410
0.400
0.407
0.380
0.389
0.426
0.441
0.459
0.444
0.553
0.496
336
0.414
0.404
0.426
0.420
0.424
0.415
0.399 0.410
0.532
0.515
0.428
0.425
0.410
0.411
0.435
0.428
0.438
0.438
0.413
0.413
0.445
0.459
0.495
0.464
0.621
0.537
720
0.475
0.438
0.491
0.459
0.487
0.450
0.454 0.439
0.666
0.589
0.487
0.461
0.478
0.450
0.499
0.462
0.527
0.502
0.474
0.453
0.543
0.490
0.585
0.516
0.671
0.561
Avg
0.403
0.398
0.407
0.410
0.414
0.407
0.387 0.400
0.513
0.496
0.419
0.419
0.400
0.406
0.429
0.425
0.435
0.437
0.403
0.407
0.448
0.452
0.481
0.456
0.588
0.517
ETTm2
96
0.181
0.264 0.180 0.264
0.182
0.265
0.175 0.259
0.287
0.366
0.207
0.305
0.187
0.267
0.189
0.280
0.209
0.308
0.193
0.292
0.203
0.287
0.192
0.274
0.255
0.339
192
0.246 0.304
0.250
0.309
0.246 0.304 0.241 0.302
0.414
0.492
0.290
0.364
0.249
0.309
0.253
0.319
0.311
0.382
0.284
0.362
0.269
0.328
0.280
0.339
0.281
0.340
336
0.307 0.343
0.311
0.348
0.307 0.342 0.305 0.343
0.597
0.542
0.377
0.422
0.321
0.351
0.314
0.357
0.442
0.466
0.369
0.427
0.325
0.366
0.334
0.361
0.339
0.372
720
0.407 0.398
0.412
0.407
0.407 0.398 0.402 0.400
1.730
1.042
0.558
0.524
0.408
0.403
0.414
0.413
0.675
0.587
0.554
0.522
0.421
0.415
0.417
0.413
0.433
0.432
Avg
0.286 0.327
0.288
0.332
0.286 0.327 0.281 0.326
0.757
0.610
0.358
0.404
0.291
0.333
0.293
0.342
0.409
0.436
0.350
0.401
0.305
0.349
0.306
0.347
0.327
0.371
ETTh1
96
0.383 0.391
0.386
0.405
0.386
0.395
0.414
0.419
0.423
0.448
0.479
0.464
0.384
0.402
0.494
0.479
0.424
0.432
0.386
0.400
0.376
0.419
0.513
0.491
0.449
0.459
192
0.433 0.419
0.441
0.436
0.437
0.424
0.460
0.445
0.471
0.474
0.525
0.492
0.436
0.429
0.538
0.504
0.475
0.462
0.437
0.432
0.420
0.448
0.534
0.504
0.500
0.482
336
0.475 0.441
0.487
0.458
0.479
0.446
0.501
0.466
0.570
0.546
0.565
0.515
0.491
0.469
0.574
0.521
0.518
0.488
0.481
0.459
0.459
0.465
0.588
0.535
0.521
0.496
720
0.475 0.464
0.503
0.491
0.481 0.470
0.500
0.488
0.653
0.621
0.594
0.558
0.521
0.500
0.562
0.535
0.547
0.533
0.519
0.516
0.506
0.507
0.643
0.616
0.514
0.512
Avg
0.441 0.429
0.454
0.447
0.446
0.434
0.469
0.454
0.529
0.522
0.541
0.507
0.458
0.450
0.542
0.510
0.491
0.479
0.456
0.452
0.440
0.460
0.570
0.537
0.496
0.487
ETTh2
96
0.288 0.337 0.297
0.349
0.288 0.338
0.302
0.348
0.745
0.584
0.400
0.440
0.340
0.374
0.340
0.391
0.397
0.437
0.333
0.387
0.358
0.397
0.476
0.458
0.346
0.388
192
0.375 0.392
0.380
0.400
0.374 0.390
0.388
0.400
0.877
0.656
0.528
0.509
0.402
0.414
0.430
0.439
0.520
0.504
0.477
0.476
0.429
0.439
0.512
0.493
0.456
0.452
336
0.423
0.434
0.428
0.432 0.415 0.426
0.426
0.433
1.043
0.731
0.643
0.571
0.452
0.452
0.485
0.479
0.626
0.559
0.594
0.541
0.496
0.487
0.552
0.551
0.482
0.486
720
0.434
0.446
0.427 0.445 0.420 0.440
0.431
0.446
1.104
0.763
0.874
0.679
0.462
0.468
0.500
0.497
0.863
0.672
0.831
0.657
0.463
0.474
0.562
0.560
0.515
0.511
Avg
0.380 0.402
0.383
0.407
0.374 0.398
0.387
0.407
0.942
0.684
0.611
0.550
0.414
0.427
0.439
0.452
0.602
0.543
0.559
0.515
0.437
0.449
0.526
0.516
0.450
0.459
Electricity
96
0.197
0.273
0.148 0.240
0.201
0.281
0.181
0.270
0.219
0.314
0.237
0.329
0.168
0.272
0.187
0.304
0.207
0.307
0.197
0.282
0.193
0.308
0.169
0.273
0.201
0.317
192
0.197
0.276
0.162 0.253
0.201
0.283
0.188
0.274
0.231
0.322
0.236
0.330
0.184
0.289
0.199
0.315
0.213
0.316
0.196
0.285
0.201
0.315
0.182
0.286
0.222
0.334
336
0.212
0.292 0.178 0.269
0.215
0.298
0.204
0.293
0.246
0.337
0.249
0.344
0.198
0.300
0.212
0.329
0.230
0.333
0.209
0.301
0.214
0.329
0.200
0.304
0.231
0.338
720
0.253
0.325
0.225
0.317
0.257
0.331
0.246
0.324
0.280
0.363
0.284
0.373
0.220 0.320
0.233
0.345
0.265
0.360
0.245
0.333
0.246
0.355
0.222
0.321
0.254
0.361
Avg
0.215
0.292
0.178 0.270
0.219
0.298
0.205
0.290
0.244
0.334
0.251
0.344
0.192
0.295
0.208
0.323
0.229
0.329
0.212
0.300
0.214
0.327
0.193
0.296
0.227
0.338
Exchange
96
0.082 0.198
0.086
0.206
0.093
0.217
0.088
0.205
0.256
0.367
0.094
0.218
0.107
0.234
0.085 0.204
0.116
0.262
0.088
0.218
0.148
0.278
0.111
0.237
0.197
0.323
192
0.176 0.297 0.177 0.299
0.184
0.307
0.176 0.299
0.470
0.509
0.184
0.307
0.226
0.344
0.182
0.303
0.215
0.359
0.176
0.315
0.271
0.380
0.219
0.335
0.300
0.369
336
0.328
0.412
0.331
0.417
0.351
0.432
0.301 0.397
1.268
0.883
0.349
0.431
0.367
0.448
0.348
0.428
0.377
0.466
0.313
0.427
0.460
0.500
0.421
0.476
0.509
0.524
720
0.861
0.700
0.847
0.691
0.886
0.714
0.901
0.714
1.767
1.068
0.852
0.698
0.964
0.746
1.025
0.774
0.831
0.699
0.839 0.695
1.195
0.841
1.092
0.769
1.447
0.941
Avg
0.362
0.402 0.360 0.403
0.378
0.417
0.367
0.404
0.940
0.707
0.370
0.413
0.416
0.443
0.410
0.427
0.385
0.447
0.354
0.414
0.519
0.500
0.461
0.454
0.613
0.539
Traffic
96
0.645
0.383
0.395 0.268
0.649
0.389
0.462
0.295
0.522
0.290
0.805
0.493
0.593
0.321
0.607
0.392
0.615
0.391
0.650
0.396
0.587
0.366
0.612
0.338
0.613
0.388
192
0.598
0.360
0.417 0.276
0.601
0.366
0.466
0.296
0.530
0.293
0.756
0.474
0.617
0.336
0.621
0.399
0.601
0.382
0.598
0.370
0.604
0.373
0.613
0.340
0.616
0.382
336
0.605
0.362
0.433 0.283
0.609
0.369
0.482 0.304
0.558
0.305
0.762
0.477
0.629
0.336
0.622
0.396
0.613
0.386
0.605
0.373
0.621
0.383
0.618
0.328
0.622
0.337
720
0.644
0.382
0.467 0.302
0.647
0.387
0.514 0.322
0.589
0.328
0.719
0.449
0.640
0.350
0.632
0.396
0.658
0.407
0.645
0.394
0.626
0.382
0.653
0.355
0.660
0.408
Avg
0.623
0.372
0.428 0.282
0.626
0.378
0.481 0.304
0.550
0.304
0.760
0.473
0.620
0.336
0.621
0.396
0.622
0.392
0.625
0.383
0.610
0.376
0.624
0.340
0.628
0.379
Weather
96
0.182
0.227
0.174
0.214
0.192
0.232
0.177
0.218 0.158
0.230
0.202
0.261
0.172
0.220
0.197
0.281
0.182
0.242
0.196
0.255
0.217
0.296
0.173
0.223
0.266
0.336
192
0.227
0.262
0.221
0.254
0.240
0.271
0.225
0.259 0.206
0.277
0.242
0.298
0.219
0.261
0.237
0.312
0.227
0.287
0.237
0.296
0.276
0.336
0.245
0.285
0.307
0.367
336
0.279
0.298
0.278 0.296
0.292
0.307
0.278 0.297 0.272
0.335
0.287
0.335
0.280
0.306
0.298
0.353
0.282
0.334
0.283
0.335
0.339
0.380
0.321
0.338
0.359
0.395
720
0.352
0.344
0.358
0.347
0.364
0.353
0.354
0.348
0.398
0.418
0.351
0.386
0.365
0.359
0.352
0.288
0.352
0.386
0.345
0.381
0.403
0.428
0.414
0.410
0.419
0.428
Avg
0.260
0.283
0.258 0.278
0.272
0.291
0.259 0.281 0.259
0.315
0.271
0.320
0.259
0.287
0.271
0.334
0.261
0.312
0.265
0.317
0.309
0.360
0.288
0.314
0.338
0.382
$1^{\text{st}}$ Count
4
15 10 15
5
6 12
5
3
0
0
0
1
0
0
1
1
0
3
0
4
0
0
0
0
0
$2^{\text{nd}}$ Count 11 9
7
6
5
8
7
17
1
3
1
0
9
1
1
1
0
0
3
1
0
0
2
0
0
0
4.2 Comparative Baseline
To thoroughly evaluate the performance of the Time Evidence Fusion Network (TEFN), we compare it against a comprehensive set of state-of-the-art baseline models. These models represent various approaches and techniques commonly used in time series forecasting, providing a robust benchmark for assessing TEFN’s capabilities. The selected baseline models are as follows:
  • Transformer-based Models:
    • iTransformer (ICLR 2024) [48]: A novel transformer architecture specifically designed for time series forecasting, showcasing the potential of transformer models in capturing long-range dependencies and complex temporal patterns.
    • PatchTST (ICLR 2023) [36]: This model divides the time series into patches and applies transformer-based operations on each patch independently. It demonstrates the effectiveness of patch-based approaches in reducing the computational complexity of transformer models while maintaining accuracy.
    • Crossformer (ICLR 2023) [20]: This model utilizes cross-attention mechanisms to capture interactions between different dimensions of time series data, demonstrating the effectiveness of transformer-based approaches in handling multivariate time series.
    • FEDformer (ICML 2022) [52]: A frequency-enhanced decomposed transformer model that captures both low-frequency and high-frequency components of time series data.
    • ETSformer [51]: A model that combines exponential smoothing and transformer-based approaches to capture both linear and non-linear trends in time series data.
    • Stationary (NeurIPS 2022) [42]: A model that utilizes a non-stationary transformer architecture to capture the evolving nature of time series data.
    • Autoformer (NeurIPS 2021) [45]: A self-supervised transformer model that learns to forecast future values without relying on explicit labels. It leverages self-attention mechanisms to capture complex temporal patterns and relationships within the data.
  • Linear Models:
    • Rlinear (ICLR 2024) [49]: This model uses a single linear layer with reversible normalization (RevIN) to forecast time series data. It is effective at capturing periodic patterns and works well with longer input sequences.
    • Dlinear (AAAI 2023) [39]: A linear model that employs seasonal-trend decomposition to capture the trend and seasonal components of time series data. It serves as a simple yet effective baseline for comparison.
    • LightTS [40]: A lightweight model that utilizes a sampling-oriented MLP structure to achieve efficient and accurate time series forecasting.
  • Other Models:
    • TiDE (ICLR 2023) [50]: A time series dense encoder model that efficiently captures temporal patterns using a dense connection structure.
    • TimesNet (ICLR 2023) [43]: A model that utilizes a temporal 2D-variation modeling approach to capture both time and frequency information in time series data.
By comparing TEFN against this diverse set of baseline models, we can comprehensively assess its performance and identify its unique advantages and contributions to the field of time series forecasting.
4.3 Implementation Details
Due to limited experimental equipment, our experimental environment was a regular desktop computer with a single CPU and a single GPU without paralleled computing of multi-GPU. Our experimental equipment is configured with AMD Ryzen9 7950x 4.5GHz 16 Cores, 64GB RAM DDR5 4800MHz, and NVIDIA RTX 4090 D 24GB. The proposed model TEFN is implemented in the PyTorch [53]. For the sake of fairness, we used an open-source benchmark  [^1] The open-source benchmark Time-Series-Library located at https://github.com/thuml/Time-Series-Library which contains comparison models and corresponding experiment configurations. The open-source address for our model is https://github.com/ztxtech/Time-Evidence-Fusion-Network.. The loss function selected during the experiment was Mean Square Error (MSE), and the optimizer selected Adam [54]. We used Mean Absolute Error (MAE) and MSE in Equation (15) and Equation (16) as error indicators in the experiment where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the predicted value.
$MAE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|$ (15)
$MSE(y,\hat{y})=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$ (16)
4.4 Forecasting Result
In the comparative experiment, we conducted a thorough comparison of 5 datasets and 12 comparative models in terms of forecasting length $L_{pred}\in\{96,192,336,720\}$ in Table II. The results showcase the remarkable capabilities of the Time Evidence Fusion Network (TEFN) in long-term time series forecasting. TEFN consistently achieves state-of-the-art (SOTA) performance across various datasets and prediction horizons, highlighting its effectiveness and versatility in capturing complex temporal patterns. Notably, TEFN achieves this with a significantly smaller model size compared to models such as iTransformer [48] and PatchTST [36], demonstrating its efficiency and scalability.
TEFN has significant advantages over other models in terms of prediction accuracy, as evidenced by the MSE and MAE metrics. Specifically, in terms of MSE, TEFN is $0.028$ smaller than iTransformer [48] and $0.025$ smaller than PatchTST [36] in the task of predicting a length of $L_{pred}=720$ on the ETTh1 dataset [19]. Similarly, in terms of MAE, TEFN once again showed a clear lead over iTransformer. TEFN was $0.008$ smaller than iTransformer [48] and $0.007$ smaller than PatchTST [36] in predicting a length of $L_{pred}=96$ on the Exchange dataset [45]. These results highlight the effectiveness and robustness of TEFN in time series prediction, demonstrating its ability to provide more accurate predictions compared to contemporary methods.
(a) TEFN (Ours)-Electricity
(b) PatchTST-Electricity
(c) Crossformer-Electricity
(d) TimesNet-Electricity
(e) Dlinear-Electricity
(f) Stationary-Electricity
(g) TEFN (Ours)-ETTm1
(h) PatchTST-ETTm1
(i) Crossformer-ETTm1
(j) TimesNet-ETTm1
(k) Dlinear-ETTm1
(l) Stationary-ETTm1
Figure 5: Visualization of long-term forecasting tasks: Figure 5(a) - Figure 5(f) in Electricity dataset with $L_{pred}=96$ and Figure 5(g) - Figure 5(l) in ETTm1 dataset with $L_{pred}=96$
We visualized the predictions of TEFN and comparative models to better demonstrate the predictive performance of different models. We take the task of Electricity-96 [44] and ETTm1-96 as an example and selected typical models PatchTST [36], Crossformer [20], TimesNet [43], Dlinear [39] and Stationary [42] for visualization in Figure 5. From the visualization results, TEFN fits the details and trends of time series better.
4.5 Efficiency Comparison
TEFN is a compact model compared to the parameters of the Transformer. In order to effectively compare the efficiency of the model, we chose the dataset Electricity with a large dataset for performance comparison. In the task Electricity-96, we measured the average time, MSE, and exported model file size for iterating over one sample. Different models occupy varying amounts of GPU and memory. Therefore, we used the size of the saved binary model files as the model size. We visualized this with a bubble chart in Figure 6 . It can be seen that TEFN achieves a low prediction error with an extremely small number of parameters. Meanwhile, it has the shortest average time for iterating over one sample, demonstrating its high efficiency. TEFN maintains almost Dlinear [39] performance and almost PatchTST [36] accuracy, making it a performance error balanced model.
Figure 6: Model efficiency comparison. Results are from the prediction task Electricity-96, where the model size refers to the size of the binary model files, and iteration time is averaged each iteration. The models beyond the canvas have not been drawn.
It is particularly relevant to note that TEFN has demonstrated strong performance on portable devices. As evidenced by the official code repository  [^2] https://github.com/ztxtech/Time-Evidence-Fusion-Network, TEFN’s training and prediction processes have been successfully executed on a Macbook Air 2020 equipped with Apple Silicon M1 and 8GB RAM. This finding suggests that TEFN’s computational efficiency is well-suited for resource-constrained environments, contrasting with the challenges typically encountered by Transformer-based models such as Crossformer [20] when deployed on portable computers.
According to the process of BPA, the time complexity of BPA is $O(n*2^{|S|})$ where $n$ is input size and $|S|$ is sample space size. For the TEFN model, the specific time complexity is $O(C*L*2^{|S|})$ where $C$ is channel size, and $L$ is the length of time series. $C$ and $|S|$ are constants and satisfy $C,|S|<<L$. Hence, the final time complexity is $O(L)$. So TEFN can maintain positive performance in long-term forecasting tasks. Another factor affecting experimental efficiency is whether TEFN is linear or nonlinear. We chose to use a linear BPA module in the experimental section because of its high efficiency. Nonlinear modules also exist, and only the membership functions need to be replaced. BPA needs to correspond to different types of nonlinear time series and design membership functions. In the future, we will explore adaptive nonlinear functions. The fitting of nonlinear functions also brings higher computational costs, therefore, in the case of efficiency priority, the linear version of TEFN is the preferred choice.
4.6 Hyperparameter Sensitivity
(a) $lr$ $|S|$MSE
(b) Violin diagram of $|S|$ and MSE
Figure 7: Traversing error of Electricity 96 task: 3D visualization and violin diagram. The points on the number axis of the violin chart reflect the density of learning rate.
To test the hyperparameter sensitivity of TEFN, we conducted a 2D traversal of the learning rate $lr\in\{10^{-2},5*10^{-2},10^{-1}\}$ and sample space size $|S|\in\{0,1,2,3,4,5,6\}$. We take the Electricity-96 task as an example. TEFN demonstrates weak sensitivity to parameters, but as the sample space increases, the error magnitude also increases, leading to a certain degree of overfitting in the model as shown in Figure 7. It is worth noting that the order of magnitude of the MSE in Figure 7 is $10^{-4}$, while the order of magnitude of the compared MSE is $10^{-3}$, so the fluctuation of different hyperparameter in error metrics is small. In order to more intuitively reflect the hyperparameter sensitivity of TEFN, correlation analysis [55, 56]  [^3] The implementation of relevant calculations for discrete indicators such as Dataset comes from dython Package (https://shakedzy.xyz/dython/). is conducted on prediction length (Target), dataset, sample space (S), Learning rate (Lr), MAE and MSE in Figure 8. According to the correlation between different indicators, it can be found that TEFN is stable to the learning rate and sample space size which are almost uncorrelated with the two error indicators MSE and MAE.
Additionally, we conducted statistical analysis on the error variance of different tasks as shown in Table III. The order of magnitude of variance is $10^{-4}$, indicating extremely small fluctuation. Random parameter selection within the traversal range of TEFN yields stable prediction results.
TABLE III: The variance of model error for traversal learning rate and sample space
Metric Dataset
($Var$) Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 0.0000 0.0000 0.0002 0.0002 0.0001 0.0000 0.0000 0.0001
192 0.0001 0.0000 0.0003 0.0004 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0002 0.0004 0.0000 0.0023 0.0000 0.0001
720 0.0000 0.0000 0.0004 0.0001 0.0000 0.0038 0.0000 0.0001
MAE 96 0.0000 0.0000 0.0001 0.0001 0.0001 0.0000 0.0000 0.0002
192 0.0001 0.0000 0.0001 0.0001 0.0000 0.0000 0.0000 0.0002
336 0.0001 0.0000 0.0001 0.0001 0.0000 0.0009 0.0000 0.0001
720 0.0000 0.0000 0.0001 0.0000 0.0000 0.0007 0.0000 0.0000
Figure 8: Heatmap of hyperparameter sensitivity correlation
4.7 Ablation Study
TEFN is formed by the fusion of time information sources $T$ and channel information sources $C$. In the ablation experiment, we made separate predictions for the time and channel parts of TEFN as shown in Table IV. Across most forecasting tasks, the fusion $C+T$ error is lower than that of individual time $T$ or channel $C$ error. Although partial use of a single channel may further reduce prediction error, the error reduction is not significant (the error of Electricity only varies $0.001$). It displays that TEFN effectively filters out time $T$ and channel $C$ information, and individual time $T$ and channel $C$ information alone cannot effectively construct the evolution pattern of time series.
We conducted ablation experiments on all forecasting tasks, testing individual channel $C$ forecasting and individual time $T$ forecasting. We tested the forecasting length $L_{pred}\in\{96,192,336,720\}$ and averaged $Avg$ the errors of the four forecasting lengths in Table IV. In the aspect of all the results of ablation, TEFN can effectively improve performance by combining the dimensions of time $T$ and channel $C$ in most cases. There are a few cases where the effect is not as good as the individual dimensions, but the difference is small and can be ignored. According to the characteristics of neural networks, the forecasting error of TEFN $l_{C+T}$ will be less than or equal to using separate time $l_{T}$ or channel $l_{C}$ in Equation (17).
$l_{C+T}\leqslant min(l_{C},l_{T})$ (17)
TABLE IV: The results of the ablation experiment
Metric Target Model Dataset
Electricity ETTh1 ETTh2 ETTm1 ETTm2 Exchange Traffic Weather
MSE 96 C+T 0.197 0.383 0.288 0.343 0.181 0.082 0.645 0.182
C 0.196$\downarrow$ 0.386$\uparrow$ 0.288 0.339$\downarrow$ 0.185$\uparrow$ 0.083$\uparrow$ 0.644$\downarrow$ 0.185$\uparrow$
T 0.197 0.386$\uparrow$ 0.291$\uparrow$ 0.356$\uparrow$ 0.182$\uparrow$ 0.083$\uparrow$ 0.645 0.195$\uparrow$
192 C+T 0.197 0.433 0.375 0.381 0.246 0.176 0.598 0.227
C 0.196$\downarrow$ 0.435$\uparrow$ 0.375 0.382$\uparrow$ 0.246 0.179$\uparrow$ 0.597$\downarrow$ 0.230$\uparrow$
T 0.197 0.435$\uparrow$ 0.375 0.390$\uparrow$ 0.247$\uparrow$ 0.175$\downarrow$ 0.598 0.240$\uparrow$
336 C+T 0.212 0.475 0.423 0.414 0.307 0.328 0.605 0.279
C 0.210$\downarrow$ 0.479$\uparrow$ 0.423 0.411$\downarrow$ 0.307 0.336$\uparrow$ 0.604$\downarrow$ 0.280$\uparrow$
T 0.212 0.479$\uparrow$ 0.414$\downarrow$ 0.423$\uparrow$ 0.307 0.324$\downarrow$ 0.606$\uparrow$ 0.291$\uparrow$
720 C+T 0.253 0.475 0.434 0.475 0.407 0.861 0.644 0.352
C 0.253 0.478$\uparrow$ 0.457$\uparrow$ 0.472$\downarrow$ 0.407 0.947$\uparrow$ 0.642$\downarrow$ 0.354$\uparrow$
T 0.254$\uparrow$ 0.479$\uparrow$ 0.442$\uparrow$ 0.487$\uparrow$ 0.408$\uparrow$ 0.871$\uparrow$ 0.643$\downarrow$ 0.364$\uparrow$
Avg C+T 0.215 0.441 0.380 0.403 0.286 0.362 0.623 0.260
C 0.214 $\downarrow$ 0.444 $\uparrow$ 0.386 $\uparrow$ 0.401 $\downarrow$ 0.286 0.386 $\uparrow$ 0.622 $\downarrow$ 0.262 $\uparrow$
T 0.215 0.445 $\uparrow$ 0.380 0.414 $\uparrow$ 0.286 0.363 $\uparrow$ 0.623 0.273 $\uparrow$
MAE 96 C+T 0.274 0.391 0.338 0.367 0.264 0.198 0.384 0.227
C 0.273$\downarrow$ 0.394$\uparrow$ 0.339$\uparrow$ 0.366$\downarrow$ 0.269$\uparrow$ 0.199$\uparrow$ 0.384 0.228$\uparrow$
T 0.274 0.393$\uparrow$ 0.339$\uparrow$ 0.377$\uparrow$ 0.265$\uparrow$ 0.199$\uparrow$ 0.383$\downarrow$ 0.234$\uparrow$
192 C+T 0.276 0.420 0.392 0.384 0.304 0.297 0.360 0.262
C 0.275$\downarrow$ 0.421$\uparrow$ 0.393$\uparrow$ 0.385$\uparrow$ 0.305$\uparrow$ 0.300$\uparrow$ 0.361$\uparrow$ 0.264$\uparrow$
T 0.276 0.421$\uparrow$ 0.390$\downarrow$ 0.392$\uparrow$ 0.305$\uparrow$ 0.296$\downarrow$ 0.361$\uparrow$ 0.270$\uparrow$
336 C+T 0.292 0.441 0.435 0.405 0.343 0.412 0.363 0.298
C 0.291$\downarrow$ 0.443$\uparrow$ 0.434$\downarrow$ 0.403$\downarrow$ 0.343 0.419$\uparrow$ 0.363 0.298
T 0.292 0.445$\uparrow$ 0.425$\downarrow$ 0.413$\uparrow$ 0.342$\downarrow$ 0.410$\downarrow$ 0.362$\downarrow$ 0.306$\uparrow$
720 C+T 0.325 0.465 0.446 0.438 0.399 0.700 0.382 0.344
C 0.325 0.463$\downarrow$ 0.462$\uparrow$ 0.436$\downarrow$ 0.399 0.739$\uparrow$ 0.382 0.347$\uparrow$
T 0.324$\downarrow$ 0.469$\uparrow$ 0.454$\uparrow$ 0.449$\uparrow$ 0.398$\downarrow$ 0.702$\uparrow$ 0.382 0.354$\uparrow$
Avg C+T 0.292 0.429 0.403 0.398 0.327 0.402 0.372 0.283
C 0.291 $\downarrow$ 0.430 $\uparrow$ 0.407 $\uparrow$ 0.397 $\downarrow$ 0.329 $\uparrow$ 0.414 $\uparrow$ 0.373 $\uparrow$ 0.284 $\uparrow$
T 0.292 0.432 $\uparrow$ 0.402 $\downarrow$ 0.408 $\uparrow$ 0.328 $\uparrow$ 0.402 0.372 0.291 $\uparrow$
$\uparrow$: Lager Error than C+T   $\downarrow$: Smaller Error than C+T
(a) Channel BPA
(b) Time BPA
Figure 9: Visualization of Time BPA and Channel BPA modules in target ETTh1-96: Each line represents the fuzzy set membership function corresponding to the time. The fuzzy numbers established in TEFN are all triangular fuzzy numbers, and different channels $C_{i}$ and time $T_{i}$ can correspond to fuzzy sets that conform to symbolic logic.
4.8 Robustness Experiments
TEFN is a highly robust neural network. To verify the robustness of TEFN in real prediction, a comparative experiment was conducted by adding noise to the dataset. The unit noise $\gamma$ follows the Gaussian distribution $N(0,1)$ with an expectation of 00 and a variance of $1$. Assume that the input time series is $x$, with time and channel dimensions. To adapt to the scale of the data, the transformation as Eq.(18) is applied to input data where $i$ is the index of channel dimension, $i$ is the index of time dimension, $x_{i}$ is a vector containing all time steps of channel $i$ and $std(\cdot)$ is Standard deviation function. So, the noisy data experiments are compared with the original data experiments.
$x^{\prime}_{ij}=x_{ij}+std(x_{i})*\gamma$ (18)
The experimental results of robustness testing are shown in the Table V. Continuing from the analysis, the experimental outcomes underscore the robust nature of TEFN, where the introduction of Gaussian noise to the dataset led to only marginal increases in prediction errors, with most variations being insignificant relative to the original errors. Notably, in long-term forecasting scenarios, a counterintuitive improvement in accuracy was observed in some instances, indicating that TEFN’s adaptability and learning capacity in complex, dynamic environments extend beyond mere resilience to noise, potentially leveraging subtle perturbations to enhance prediction performance.
TABLE V: The results of robustness experiments
Dataset Pred Original Noisy
MSE MAE MSE MAE
ETTm1 96 0.343 0.367 0.414 ($\uparrow$ 0.071) 0.411 ($\uparrow$ 0.044)
192 0.381 0.383 0.443 ($\uparrow$ 0.062) 0.429 ($\uparrow$ 0.046)
336 0.414 0.404 0.499 ($\uparrow$ 0.085) 0.458 ($\uparrow$ 0.054)
720 0.475 0.438 0.381 ($\downarrow$ 0.094) 0.398 ($\downarrow$ 0.040)
ETTm2 96 0.181 0.264 0.254 ($\uparrow$ 0.073) 0.312 ($\uparrow$ 0.048)
192 0.246 0.304 0.314 ($\uparrow$ 0.068) 0.350 ($\uparrow$ 0.046)
336 0.307 0.343 0.411 ($\uparrow$ 0.104) 0.403 ($\uparrow$ 0.060)
720 0.407 0.398 0.199 ($\downarrow$ 0.208) 0.281 ($\downarrow$ 0.117)
ETTh1 96 0.383 0.391 0.466 ($\uparrow$ 0.083) 0.448 ($\uparrow$ 0.057)
192 0.433 0.419 0.508 ($\uparrow$ 0.075) 0.468 ($\uparrow$ 0.049)
336 0.475 0.441 0.509 ($\uparrow$ 0.034) 0.484 ($\uparrow$ 0.043)
720 0.475 0.464 0.426 ($\downarrow$ 0.049) 0.427 ($\downarrow$ 0.037)
ETTh2 96 0.288 0.337 0.386 ($\uparrow$ 0.098) 0.401 ($\uparrow$ 0.064)
192 0.375 0.392 0.432 ($\uparrow$ 0.057) 0.441 ($\uparrow$ 0.049)
336 0.423 0.434 0.466 ($\uparrow$ 0.043) 0.467 ($\uparrow$ 0.033)
720 0.434 0.446 0.304 ($\downarrow$ 0.130) 0.352 ($\downarrow$ 0.094)
Electricity 96 0.197 0.273 0.270 ($\uparrow$ 0.073) 0.363 ($\uparrow$ 0.090)
192 0.197 0.276 0.281 ($\uparrow$ 0.084) 0.371 ($\uparrow$ 0.095)
336 0.212 0.292 0.322 ($\uparrow$ 0.110) 0.399 ($\uparrow$ 0.107)
720 0.253 0.325 0.270 ($\uparrow$ 0.017) 0.361 ($\uparrow$ 0.036)
Traffic 96 0.645 0.383 0.727 ($\uparrow$ 0.082) 0.479 ($\uparrow$ 0.096)
192 0.598 0.360 0.734 ($\uparrow$ 0.136) 0.481 ($\uparrow$ 0.121)
336 0.605 0.362 0.768 ($\uparrow$ 0.163) 0.491 ($\uparrow$ 0.129)
720 0.644 0.382 0.774 ($\uparrow$ 0.130) 0.494 ($\uparrow$ 0.112)
Weather 96 0.182 0.227 0.232 ($\uparrow$ 0.050) 0.269 ($\uparrow$ 0.042)
192 0.227 0.262 0.285 ($\uparrow$ 0.058) 0.305 ($\uparrow$ 0.043)
336 0.279 0.298 0.357 ($\uparrow$ 0.078) 0.350 ($\uparrow$ 0.052)
720 0.352 0.344 0.190 ($\downarrow$ 0.162) 0.237 ($\downarrow$ 0.107)
Exchange 96 0.082 0.198 0.188 ($\uparrow$ 0.106) 0.308 ($\uparrow$ 0.110)
192 0.176 0.297 0.340 ($\uparrow$ 0.164) 0.422 ($\uparrow$ 0.125)
336 0.328 0.412 0.963 ($\uparrow$ 0.635) 0.739 ($\uparrow$ 0.327)
720 0.861 0.700 0.092 ($\downarrow$ 0.769) 0.213 ($\downarrow$ 0.487)
4.9 Interpretability Analysis
The Time Evidence Fusion Network (TEFN) incorporates fuzzy logic and the Dempster-Shafer theory to effectively model the inherent uncertainty and ambiguity within time series data, resulting in a model that is not only accurate but also highly interpretable. TEFN allocates mass to events within the event space $E=\{E_{i}|i\in[1,2^{|S|}]\}$, where each event represents a potential outcome or scenario. The mass assigned to an event reflects the level of confidence or support for that event, providing insight into the model’s belief about its likelihood of occurrence. This support allocation is particularly interpretable due to the use of fuzzy membership functions $\mu(x)=wx+b$ within the BPA module, as depicted in Figure 9. By analyzing the shape and parameters of these membership functions, users can gain a clear understanding of the model’s perception of data features and their influence on predictions. For instance, it becomes evident which features are deemed more important than others, and how the model responds to variations within these features. This visualization not only makes the model’s inner workings more transparent but also allows users to intuitively assess the model’s confidence in its predictions and the level of uncertainty associated with each outcome.
In summary, TEFN’s interpretability stems from its foundational reliance on fuzzy logic and the Dempster-Shafer theory, coupled with the transparency of its BPA and fuzzy reasoning processes. By exploring these aspects, users can develop a comprehensive understanding of the model’s decision-making process and prediction outcomes, fostering trust and confidence in the model’s capabilities.
4.10 Limitation and Future Prospects
TEFN is a naive and unmodified model. There is not engineering experience in the application of BPA modules, such as how to generate BPA, and fuzzy logic is not the optimal implementation method. Due to fuzzy logic, TEFN is essentially a linear model that cannot handle some nonlinear time series well. In addition, for the fusion of mass distribution in the face of the failure of the classic DSR for a large amount of data, using expectations as the fusion method may not necessarily be the optimal pattern. In terms of performance optimization, TEFN adopts a simple fully connected approach, which can lead to a rapid increase in parameter count for time series that are too long or have multiple channels. In fact, the implementation techniques of TEFN can further optimize the performance of the model through sampling and kernel operations similar to convolution. Thanks to the fact that evidence theory belongs to symbolic logic, the interpretation of neural network parameters will be more in line with human logic. Effective parameter initialization may exist, for example, initializing corresponding parameters in a non-neural network way may lead to faster convergence.
5 Conclusion
This paper introduces the TEFN, a novel approach for long-term time series forecasting that is particularly well-suited for handling very large datasets. TEFN treats different channels and time points as distinct information sources, using evidence theory to construct basic probability assignments and discard irrelevant data. This selective information fusion significantly reduces the computational burden, enabling TEFN to process and predict within acceptable time frames efficiently. The model’s parameter efficiency, robustness, and interpretability make it a valuable tool for time series forecasting tasks involving large and complex datasets.